OpenStudy (anonymous):

How many mL of 1.50 M HCl are needed to titrate 25.0 mL of 1.20 M Ca(OH)2?

OpenStudy (anonymous):

Any idea ??

OpenStudy (anonymous):

@Creativegenius

OpenStudy (anonymous):

The acid should equal the base?

OpenStudy (anonymous):

yeah so the formula for it is Ca x Va = Cb x Vb

OpenStudy (anonymous):

As we have to find Va so Va = Cb x Vb / Ca

OpenStudy (anonymous):

so would it be 20?

OpenStudy (anonymous):

mL

OpenStudy (anonymous):

I didn't solve it .. :)

OpenStudy (anonymous):

yeah its come 20mL ..

OpenStudy (anonymous):

it said i was wrong it was 40mL

OpenStudy (anonymous):

wait wait I missed Something

OpenStudy (anonymous):

Balance equation Ca(OH)2 + 2 HCl = CaCl2 + 2 H2O So formula is Ca x Va = 2 x Cb x Vb 'cause two moles of Acid ... Sorry I skipped this step .. :)

OpenStudy (anonymous):

Its ok thank you ;-)

OpenStudy (anonymous):

np :)